OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (6, 3) and has a slope of 2/3 2x + 3y = 3 3x – 2y = 3 3x + 2y = 3 2x – 3y = 3

jhonyy9 (jhonyy9):

the last is right sure

jhonyy9 (jhonyy9):

because there is 3y=2x-3 and y=(2x-3)/3 so y= (2/3)x -1